# Efficient Jailbreak Mitigation Using Semantic Linear Classification in a Multi-Staged Pipeline

### v1

@@ Abstract
Prompt injection and jailbreaking attacks pose persistent security challenges to large language model (LLM)–based systems. We present an efficient and systematically evaluated defense architecture that mitigates these threats through a lightweight, multi‑stage pipeline. Its core component is a semantic filter based on text normalization, TF–IDF representations, and a Linear SVM classifier. Despite its simplicity, this module achieves 93.4% accuracy and 96.5% specificity on held‑out data, substantially reducing attack throughput while incurring negligible computational overhead.
@@
Evaluation across a curated corpus of over 30 000 labeled prompts, including benign, jailbreak, and application‑layer injections, confirms that staged, resource‑efficient defenses can robustly secure modern LLM‑driven applications.
Efficient Jailbreak Mitigation Using Semantic Linear Classification in a Multi-Staged Pipeline
Akshaj Prashanth Rao, Advait Singh, Saumya Kumaar Saksena, Dhruv Kumar Birla Institute of Technology and Science, Pilani {f20220289, f20220636}@pilani.bits-pilani.ac.in, f2013114h@alumni.bits-pilani.ac.in, dhruv.kumar@pilani.bits-pilani.ac.in
@@ 1 Introduction
A central contribution of this work is an efficient semantic defense stage based on a Linear Support Vector Machine (LSVM) classifier operating over normalized textual representations. Designed for early deployment in a defensive pipeline, this component offers strong discriminative performance under minimal computational cost, enabling rapid screening of untrusted inputs before more complex analyses are triggered. Its lightweight footprint allows the defense pipeline to scale to high-volume inference environments without incurring significant overhead. Additionally, in comparison with well established benchmarks like ShieldGemma, our SVM‑based configuration improves overall accuracy from 35.1 % to 93.4% while reducing average time‑to‑completion from $\approx$ 450s to 47s, yielding over 10× lower latency than ShieldGemma.
@@ 3.2 Pipeline configuration
Defenses are composed and parameterized using the Hydra configuration framework. Input prompts traverse the configured pipeline before reaching the target model; outputs are similarly screened to ensure consistent enforcement across the input–output boundary. This modular orchestration supports automated benchmarking and simplifies comparison across different defensive configurations.
@@ 3.4 Defense Suite Implementation
Our defense suite is implemented as a configurable multi‑stage pipeline that screens both inputs to, and outputs from, the target LLM. Each stage functions as an independent module that receives a prompt and returns an allow/block decision together with optional metadata such as confidence scores or matched features. Pipeline behavior: stage ordering, activation, and threshold parameters, is defined programmatically through Hydra, enabling reproducible experimentation and fine‑grained control over accuracy and latency trade‑offs.
The defense stages operate sequentially: inexpensive filters are placed early to intercept clear adversarial inputs, while more computationally intensive analyses are reserved for ambiguous or residual cases. Prompts passing all input checks are forwarded to the model, and generated outputs can optionally be re-evaluated by an output scanner to catch any remaining unsafe completions.
Across all evaluated configurations, the LSVM stage, comprising text normalization, TF–IDF vectorization, and linear classification, demonstrated consistently high detection effectiveness at minimal computational cost. Its efficiency and discriminative power make it a strong early gate within the broader pipeline, significantly reducing the number of prompts requiring higher‑latency evaluation.
We benchmarked pipeline performance against ShieldGemma (see Table 2). The comparatively weaker robustness of ShieldGemma under our diverse attack mix underscores the difficulty of defending against modern prompt injections and motivates the layered strategy described in the following sections. Subsequent subsections present each defense module in detail and describe how these components combine to deliver resilient and resource‑efficient protection.
@@ 3.7 Injection-Regex
3.8 YARA pattern scanner
@@ 3.9 Output scanner
3.10 LLM-as-a-judge (baseline)
The final component is an LLM-based judge used primarily as a standardized moderation baseline. We use ShieldGemma-2B as specified in Table 1: it consumes the prompt alongside structured safety guidelines and returns a binary verdict (violation or not), optionally with a rationale. This stage is valuable for high-fidelity policy assessment but is typically higher-latency than lightweight filters; accordingly, it is positioned as a late-stage or benchmarking component rather than the primary production gate.
@@ 4.2 Experimental Setup
Table 3: Defense effectiveness on injection prompts. Prompts that passed the defense stack were forwarded to gpt-oss:20b for execution and the results of the attempts were independently verified using Gemini.
@@ 5.4 Summary
Taken together, our results demonstrate that a carefully staged, multi-layer defense pipeline can eliminate successful jailbreak and prompt-injection attacks while preserving high accuracy on benign inputs and maintaining low computational overhead. In particular, the semantic LSVM module emerges as a key contributor, offering strong generalization at minimal cost and enabling robust end-to-end protection when combined with complementary defenses.
@@ 6 Conclusion
In this work, we presented a systematic and reproducible evaluation of defenses against prompt injection and jailbreak attacks in large language model (LLM) systems. Motivated by the practical reality that no single mitigation is sufficient against rapidly evolving adversarial prompts, we framed defense as a configurable, multi-stage pipeline and evaluated its components under realistic attack pressure using a large, curated corpus of over 30,000 labeled prompts.
Our results show that lightweight, interpretable defenses can play a decisive role in securing LLM-powered applications. In particular, the proposed text processing and semantic Linear SVM (LSVM) module emerged as a highly effective early-stage filter, achieving high accuracy and specificity while imposing minimal computational overhead. When combined with complementary defenses such as vector similarity matching and sequence-based classifiers, the full pipeline successfully reduced the Attack Success Rate (ASR) to zero across all evaluated adversarial attempts, while preserving strong performance on benign prompts.
Beyond individual performance gains, our study highlights several broader insights. First, accuracy and false-positive control are critical metrics for real-world deployment, as overly conservative defenses can significantly degrade usability. Second, staged guardrail architectures enable favorable security–latency trade-offs by allowing inexpensive filters to absorb most attack traffic before invoking higher-cost mechanisms. Finally, our findings suggest that classical machine learning techniques, when carefully engineered and evaluated, remain highly competitive alongside modern LLM-based moderation approaches.
We release our dataset, implementation, and experimental framework to support reproducibility and future research. This work encourages further exploration of transparent, modular, and cost-efficient defenses, and serves as a practical reference for practitioners seeking to harden LLM systems against prompt-based attacks in real-world settings.

### v2

# PromptScreen: Efficient Jailbreak Mitigation Using Semantic Linear Classification in a Multi-Staged Pipeline

@@ Abstract
Prompt injection and jailbreaking attacks pose persistent security challenges to large language model (LLM)–based systems. We present PromptScreen, an efficient and systematically evaluated defense architecture that mitigates these threats through a lightweight, multi‑stage pipeline. Its core component is a semantic filter based on text normalization, TF–IDF representations, and a Linear SVM classifier. Despite its simplicity, this module achieves 93.4% accuracy and 96.5% specificity on held‑out data, substantially reducing attack throughput while incurring negligible computational overhead.
@@
Evaluation across a curated corpus of over 30,000 labeled prompts, including benign, jailbreak, and application‑layer injections, confirms that staged, resource‑efficient defenses can robustly secure modern LLM‑driven applications.
PromptScreen: Efficient Jailbreak Mitigation Using Semantic Linear Classification in a Multi-Staged Pipeline
Akshaj Prashanth Rao1, Advait Singh1, Saumya Kumaar Saksena1, Dhruv Kumar1,2 1Birla Institute of Technology and Science, Pilani 2Trustwise {f20220289, f20220636}@pilani.bits-pilani.ac.in, f2013114h@alumni.bits-pilani.ac.in, dhruv.kumar@pilani.bits-pilani.ac.in
@@ 1 Introduction
A central contribution of this work is an efficient semantic defense stage based on a Linear Support Vector Machine (LSVM) classifier operating over normalized textual representations. Designed for early deployment in a defensive pipeline, this component offers strong discriminative performance under minimal computational cost, enabling rapid screening of untrusted inputs before more complex analyses are triggered. Its lightweight footprint allows the defense pipeline to scale to high-volume inference environments without incurring significant overhead. Additionally, in comparison with well established benchmarks like ShieldGemma(Google, 2024), our SVM‑based configuration improves overall accuracy from 35.1 % to 93.4% while reducing average time‑to‑completion from $\approx$ 450s to 47s, yielding over 10× lower latency than ShieldGemma.
@@ 3.3 Defense Suite Implementation
The defense stages operate sequentially: inexpensive filters are placed early to intercept clear adversarial inputs, while more computationally intensive analyses are reserved for ambiguous or residual cases. Prompts passing all input checks are forwarded to the model, and generated outputs can optionally be re-evaluated by an output scanner to catch any remaining unsafe completions. Pipeline behavior: stage ordering, activation, and threshold parameters, is defined programmatically through Hydra (Mehrabi et al., 2020), enabling reproducible experimentation and fine‑grained control over accuracy and latency trade‑offs.
@@ 3.3.4 Injection-Regex
3.3.5 YARA pattern scanner
@@ 3.3.6 Output scanner
3.3.7 LLM-as-a-judge (baseline)
The final component is an LLM-based judge used primarily as a standardized moderation baseline. We use ShieldGemma-2B: it consumes the prompt alongside structured safety guidelines and returns a binary verdict (violation or not), optionally with a rationale. This stage is valuable for high-fidelity policy assessment but is typically higher-latency than lightweight filters; accordingly, it is positioned as a late-stage or benchmarking component rather than the primary production gate.
@@ 4.2 Experimental Setup
Table 3: Defense effectiveness on injection prompts. Prompts that passed the defense stack were forwarded to gpt-oss:20b for execution and the results were independently verified using Gemini.
@@ 5.1 Classification Performance
SVM Configuration Precision Sensitivity Specificity NPV Accuracy
Baseline 0.8255 0.8942 0.9069 0.9457 0.9027
Word n-gram (1,2) 0.8213 0.8914 0.9045 0.9442 0.9002
Word n-gram (1,3) 0.8218 0.8906 0.9049 0.9438 0.9002
Word bigram 0.5940 0.9261 0.6884 0.9498 0.7668
Char n-gram (2,4) 0.8895 0.9371 0.9427 0.9682 0.9409
Char n-gram (3,5) 0.8827 0.9314 0.9391 0.9653 0.9366
Hybrid (word+char) 0.8736 0.9285 0.9339 0.9637 0.9321
Hybrid (extended) 0.8748 0.9298 0.9345 0.9643 0.9329
Table 4: Performance metrics for different SVM configurations.
@@ 5.4 Summary
Overall, our results show that a staged, multi-layer defense pipeline can eliminate jailbreak and prompt-injection attacks while maintaining high accuracy on benign inputs with low computational overhead. The semantic LSVM module is central to this success, providing strong generalization at minimal cost and enabling robust end-to-end protection when combined with complementary defenses.
6 SVM Ablation Study
To better understand the factors driving the strong performance of the semantic LSVM defense, we conduct a systematic ablation study over alternative feature representations. The goals of this study are twofold: (i) to quantify the contribution of different n-gram granularities to detection performance, and (ii) to examine robustness–efficiency trade-offs across word-level, character-level, and hybrid representations. All ablated models share the same preprocessing pipeline, training procedure, and evaluation protocol, differing only in the TF–IDF feature configuration supplied to the Linear SVM.
Experimental setup.
To stress-test robustness under adversarial conditions, we construct an augmented variant of the dataset in which each original prompt is expanded into four perturbed versions, yielding over 120,000 prompts in total. Perturbations include leetspeak substitutions, Unicode homoglyphs, and whitespace manipulations applied at varying levels of intensity. Each SVM configuration is trained on the same training split described in Section 3 and evaluated on an identical held-out test set to ensure fair comparison. As in the rest of the paper, the positive class corresponds to benign prompts. We report precision, sensitivity (recall), specificity, negative predictive value (NPV), and overall accuracy.
Ablated configurations.
We evaluate four classes of feature representations:
(i) a baseline word-level TF–IDF model using unigrams,
(ii) extended word n-gram models capturing short-range compositional patterns,
(iii) character n-gram models designed to capture obfuscation, misspellings, and token fragmentation,
and (iv) hybrid models that concatenate word- and character-level features.
Results and analysis.
Table 4 summarizes the performance of all evaluated configurations. Several trends are immediately apparent. Word-level models provide a reasonable baseline but exhibit early saturation in accuracy and specificity, indicating limited robustness to paraphrasing and token-level perturbations. In particular, the word-bigram-only configuration shows a marked decline in precision and specificity, suggesting overfitting to frequent co-occurrence patterns that fail to generalize to unseen adversarial inputs.
Character n-gram models, by contrast, demonstrate substantially improved robustness and achieve the strongest overall performance. The char n-gram (2,4) configuration yields the highest accuracy at 94.09%, along with strong specificity and the highest NPV, indicating reliable rejection of adversarial prompts while preserving benign inputs. This behavior aligns with the observation that many jailbreak and prompt injection attacks rely on subtle formatting changes, token splitting, or character-level perturbations that are poorly captured by word-based representations.
Hybrid word–character representations perform competitively but do not surpass the best pure character-based configuration. While hybrids offer balanced performance, they introduce additional feature dimensionality without a commensurate improvement in accuracy, making them less attractive for latency- and memory-constrained deployments.
@@ 7 Conclusion
In this work, we presented a systematic and reproducible evaluation of defenses against prompt injection and jailbreak attacks in large language model (LLM) systems, framing mitigation as a configurable multi-stage pipeline evaluated under realistic attack pressure using a curated corpus of over 30,000 labeled prompts. Our results demonstrate that lightweight and interpretable defenses can play a decisive role in securing LLM-powered applications: in particular, the proposed text processing and semantic Linear SVM (LSVM) module emerged as an effective early-stage filter, achieving high accuracy and specificity with minimal computational overhead. When combined with complementary mechanisms such as vector similarity matching and sequence-based classifiers, the full defense stack reduced the Attack Success Rate (ASR) to zero across all evaluated adversarial attempts while preserving strong performance on benign prompts. Beyond individual performance gains, our findings highlight the importance of false-positive control for usability, the benefits of staged architectures for favorable security–latency trade-offs, and the continued competitiveness of carefully engineered classical machine learning techniques alongside modern LLM-based moderation approaches. To support reproducibility and future research, we release our dataset, implementation, and experimental framework; the source code is available at https://github.com/dronefreak/PromptScreen
8 Limitations
  1. 1.
    Multilingual Constraints: A primary limitation is that the current preprocessing pipeline is optimized for English-language prompts. Consequently, the system currently does not address or evaluate protection against multilingual jailbreak attempts.
  2. 2.
    Contextual and Multi-turn Awareness: The pipeline primarily evaluates prompts as isolated, single-turn inputs. This design may be less effective against sophisticated multi-turn attacks where adversarial intent is distributed incrementally across a conversation history rather than contained within a single message.
  3. 3.
    Generalization to Zero-Day Attacks: While the SVM shows superior generalization compared to similarity-based lookups, its performance is fundamentally bounded by the diversity of its training corpus. Truly novel “zero-day” attack strategies that utilize entirely different linguistic structures than those in the training dataset may initially bypass the semantic gate, although subsequent stages may detect adversarial intent.